For Problems 5-7, use f(x) = x ^ 3 - 3 and g(x) = 2 ^ x
5. Find (f+g)(x). 6
6. Evaluate f(4) and g(4).
7. Evaluate (f + g)(4)

Answer by math_tutor2020(3828)


Problem 5

(f+g)(x) = f(x) + g(x)
(f+g)(x) = x^3-3 + 2^x
(f+g)(x) = 2^x + x^3 - 3
There's not much to do here in terms of simplification.

---------------------------------------------------------------------
Problem 6

f(x) = x^3-3
f(4) = 4^3-3
f(4) = 64-3
f(4) = 61

g(x) = 2^x
g(4) = 2^4
g(4) = 16

---------------------------------------------------------------------
Problem 7

(f+g)(x) = 2^x + x^3 - 3
(f+g)(4) = 2^4 + 4^3 - 3
(f+g)(4) = 16 + 64 - 3
(f+g)(4) = 77

Another approach.
(f+g)(x) = f(x) + g(x)
(f+g)(4) = f(4) + g(4)
(f+g)(4) = 61 + 16
(f+g)(4) = 77